# Resona: Improving Context Copying in Linear Recurrence Models with Retrieval

### v1

@@ C.3 Training Details for Synthetic Benchmarks
We evaluate Resona on the MQAR task by training six base architectures: Base Con, H3, Hyena, Mamba, Mamba2, and RWKV. All models are trained with a 4-layer configuration and a hidden dimension (d_model) ranging from 32 to 256. The sequence lengths vary from 64 to 512, with the number of key-value (KV) pairs corresponding to 4–32, respectively. To integrate Resona, we insert the Resona Layer into either the first or third layer of each model, using the same d_model settings. The learning rate is swept using the default settings from Arora et al. [2024]. For the Resona Layer, we use a chunk size of 2 and a top-$k$ value of 1. The exact results for each model, sequence length, and hidden dimension can be found in Table 9.
Model Length=64 Length=128 Length=256 Length=512
@@
BaseConv$+$Resona0 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
BaseConv$+$Resona2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
@@
Hyena$+$Resona0 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Hyena$+$Resona2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
@@
RWKV$+$Resona0 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
RWKV$+$Resona2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
@@
Mamba$+$Resona0 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Mamba$+$Resona2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
@@
Mamba2$+$Resona0 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Mamba2$+$Resona2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
@@
C.4 Training Details for Real-World Tasks
Pretraining
During pre-training, we prioritized maintaining consistent model depth across architectures. For GLA, DeltaNet, and RetNet, we adopted the architecture implementations from FLA-Hub [Yang and Zhang, 2024], configuring them with 24 layers and a hidden size of 600. For Hymba, we utilized NVIDIA’s official 150M implementation (24 layers with a hidden size of 512). For Mamba, we employed FLA-Hub’s implementation with 48 layers and a hidden size of 600. For comparative models with equivalent parameter counts, we adjusted the hidden size from 600 to 640. Following the methodology outlined in the Hymba paper, we integrated Resonamodules at the shallowest, middle, and deepest layers to reinforce critical information flow. Each model was trained for 8,000 steps, with model selection performed using a dedicated validation set. The training configuration employed a cosine annealing scheduler with 5% warmup ratio, the AdamW optimizer (learning rate = 1e-3, weight decay = 0.01), and gradient clipping at 1.0.
Finetuning
We employed a base learning rate of 1e-5 for fine-tuning on the three individual datasets, with other training configurations remaining similar to those used in pre-training, except that Resonamodules benefited from a higher learning rate. For CoQA, NarrativeQA, and TriviaQA, we trained for 2k, 8k, and 10k steps, respectively. During general supervised fine-tuning, we created a unified training set by shuffling 10k samples from each of the three QA datasets, ensuring significant diversity in sequence length and content. Additionally, we constructed a general test set by selecting 500 validation samples from each dataset. We trained on this combined dataset for 4k steps.
C.5 NIAH Scoring Details

### v3

@@ A.3 Training Details for Synthetic Benchmarks
We evaluate Resona on the MQAR task by training six base architectures: BaseConv, H3, Hyena, Mamba, Mamba2, and RWKV. All models are trained with a 4-layer configuration and a hidden dimension (d_model) ranging from 32 to 256. The sequence lengths vary from 64 to 512, with the number of key-value (KV) pairs corresponding to 4–32, respectively. To integrate Resona, we insert the Resona Layer into either the first or third layer of each model, using the same d_model settings. The learning rate is swept using the default settings from Arora et al. [2024]. For the Resona Layer, we use a chunk size of 2 and a top-$k$ value of 1. The exact results for each model, sequence length, and hidden dimension can be found in Table 11.
Model $L=64$ $L=128$ $L=256$ $L=512$
@@
+Resona0 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+Resona2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
@@
+Resona0 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+Resona2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
@@
+Resona0 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+Resona2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
@@
+Resona0 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+Resona2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
@@
+Resona0 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+Resona2 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
@@
A.4 Training Details for Real-World Tasks
Pretraining.
During pre-training, we prioritized maintaining consistent model depth across architectures. For GLA, DeltaNet, and RetNet, we adopted the architecture implementations from Dao et al. [2022], configuring them with 24 layers and a hidden size of 600. For Hymba, we use NVIDIA’s official 150M parameter implementation (24 layers with a hidden size of 512). For Mamba, we employed FLA-Hub’s implementation with 48 layers and a hidden size of 600. For comparative models with equivalent parameter counts, we adjusted the hidden size from 600 to 640. Following the methodology outlined in the Hymba paper, we integrated Resonamodules at the shallowest, middle, and deepest layers to reinforce critical information flow. Each model was trained for 8,000 steps, with model selection performed using a dedicated validation set. The training configuration employed a cosine annealing scheduler with warmup over 5% of the training steps, the AdamW optimizer (learning rate of $\mathsf{1e\text{-}3}$ and weight decay = $\mathsf{0.01}$), and gradient clipping of $\mathsf{1.0}$.
Finetuning.
We employed a base learning rate of $\mathsf{1e\text{-}5}$ for fine-tuning on the three individual datasets, with other training configurations remaining similar to those used in pre-training, except that Resonamodules benefited from a higher learning rate. For CoQA, NarrativeQA, and TriviaQA, we trained for 2K, 8K, and 10K steps, respectively. During general supervised fine-tuning, we created a unified training set by shuffling 10K samples from each of the three QA datasets, ensuring significant diversity in sequence length and content. Additionally, we constructed a general test set by selecting 500 validation samples from each dataset. We trained on this combined dataset for 4K steps.
A.5 NIAH Scoring Details